# EP-GAT: Energy-based Parallel Graph Attention Neural Network for Stock Trend Classification

### v1

@@ Abstract
Graph neural networks have shown remarkable performance in forecasting stock movements, which arises from learning complex inter-dependencies between stocks and intra-dynamics of stocks. Existing approaches based on graph neural networks typically rely on static or manually defined factors to model changing inter-dependencies between stocks. Furthermore, these works often struggle to preserve hierarchical features within stocks. To bridge these gaps, this work presents the Energy-based Parallel Graph Attention Neural Network, a novel approach for predicting future movements for multiple stocks. First, it generates a dynamic stock graph with the energy difference between stocks and Boltzmann distribution, capturing evolving inter-dependencies between stocks. Then, a parallel graph attention mechanism is proposed to preserve the hierarchical intra-stock dynamics. Extensive experiments on five real-world datasets are conducted to validate the proposed approach, spanning from the US stock markets (NASDAQ, NYSE, SP) and UK stock markets (FTSE, LSE). The experimental results demonstrate that EP-GAT consistently outperforms competitive five baselines on test periods across various metrics. The ablation studies and hyperparameter sensitivity analysis further validate the effectiveness of each module in the proposed method.
Index Terms:
stock trend forecasting, graph neural network, dynamic graph modelling, graph attention, Boltzmann distribution.
@@ III-A Notations
Here, $\Omega_{L}(\cdot)$ denotes the model EP-GAT with $L$ propagation layers, $\mathbf{C}_{t}\in\mathbb{Z}^{N\times\phi\alpha}$ denotes the label matrix, $\phi$ denotes the forecast step, and $\alpha$ denotes the number of trends. Concretely, the model utilizes historical stock indicators from interval $[t-\tau+1,t]$ to predict future trends of multiple stocks over interval $[t+1,t+\phi]$.
@@ IV-B Parallel Graph Attention
Here, $\mathbf{H}^{\prime}_{l}$ denotes the learned node representation from the parallel graph attention mechanism, $\mathbf{H}_{l}$ denotes the intermediate node representation from propagation, ||||| | denotes the concatenation operation over the column, $\eta(\cdot)$ denotes the GATv2Conv [49] propagation operation, and $\gamma(\cdot)$ denotes the multi-head attention. The multi-head attention is expressed as,
@@ V Experiment
ACC(%) MCC(×$\times 10^{-2}$) F1-Score ACC(%) MCC(×$\times 10^{-2}$) F1-Score ACC(%) MCC(×$\times 10^{-2}$) F1-Score ACC(%) MCC(×$\times 10^{-2}$) F1-Score ACC(%) MCC(×$\times 10^{-2}$) F1-Score
GraphWaveNet [56] $52.67\pm 0.17$ $-0.73\pm 0.02$ $0.53\pm 0.02$ $53.42\pm 0.23$ $-2.57\pm 0.01$ $0.54\pm 0.01$ $52.91\pm 0.22$ $-2.39\pm 0.01$ $0.52\pm 0.01$ $53.09\pm 0.14$ $1.77\pm 0.02$ $0.53\pm 0.01$ $53.17\pm 0.19$ $-3.44\pm 0.02$ $0.52\pm 0.01$
@@ V-A2 Model Setting
In this work, the lag window size $\tau$ is searched over $[7,27]$, the constant $k$ is searched from $[0.02,2]$, and the sparsification threshold $s$ is searched from $[0.25,0.85]$. The number of stock indicators $F$ is set to 4. The number of heads $h$ is searched over $[2,30]$, and the number of layers $L$ is searched from $[2,6]$. In addition, we adopt the PReLU as the activation function, and the AdamW as the optimizer. The learning rate is searched from $[10^{-4},2\times 10^{-3}]$ weight decay is searched from $[10^{-4},10^{-3}]$, and training epoch is set to 800. In this work, the number of trends $\alpha$ is set to 2, and forecast step $\phi$ is set to 1 following previous works [26, 38, 33, 25], meaning a classification task of predicting next-day stock movements. The detailed datasets and code are available at here.

### v2

@@ Abstract
Graph neural networks have shown remarkable performance in forecasting stock movements, which arises from learning complex inter-dependencies between stocks and intra-dynamics of stocks. Existing approaches based on graph neural networks typically rely on static or manually defined factors to model changing inter-dependencies between stocks. Furthermore, these works often struggle to preserve hierarchical features within stocks. To bridge these gaps, this work presents the Energy-based Parallel Graph Attention Neural Network, a novel approach for predicting future movements for multiple stocks. First, it generates a dynamic stock graph with the energy difference between stocks and Boltzmann distribution, capturing evolving inter-dependencies between stocks. Then, a parallel graph attention mechanism is proposed to preserve the hierarchical intra-stock dynamics. Extensive experiments on five real-world datasets are conducted to validate the proposed approach, spanning from the US stock markets (NASDAQ, NYSE, SP) and UK stock markets (FTSE, LSE). The experimental results demonstrate that EP-GAT consistently outperforms competitive five baselines on test periods across various metrics. The ablation studies and hyperparameter sensitivity analysis further validate the effectiveness of each module in the proposed method. The raw dataset and code are available at https://github.com/theflash987/EP-GAT.
@@ III-A Notations
Here, $\Omega_{L}(\cdot)$ denotes the model EP-GAT with $L$ propagation layers, $\mathbf{C}_{t}\in\mathbb{Z}^{N\times\phi\alpha}$ denotes the label matrix, $\phi$ denotes the forecast step, and $\alpha$ denotes the number of trends. Concretely, the model utilizes historical stock indicators from interval $t-\tau+1,t]$ to predict future trends of multiple stocks over interval $t+1,t+\phi]$.
@@ IV-B Parallel Graph Attention
Here, $\mathbf{H}^{\prime}_{l}$ denotes the learned node representation from the parallel graph attention mechanism, $\mathbf{H}_{l}$ denotes the intermediate node representation from propagation, $||$ denotes the concatenation operation over the column, $\eta(\cdot)$ denotes the GATv2Conv [6] propagation operation, and $\gamma(\cdot)$ denotes the multi-head attention. The multi-head attention is expressed as,
@@ V Experiment
ACC(%) MCC($\times 10^{-2}$) F1-Score ACC(%) MCC($\times 10^{-2}$) F1-Score ACC(%) MCC($\times 10^{-2}$) F1-Score ACC(%) MCC($\times 10^{-2}$) F1-Score ACC(%) MCC($\times 10^{-2}$) F1-Score
GraphWaveNet [50] $52.67\pm 0.17$ $0.73\pm 0.02$ $0.53\pm 0.02$ $53.42\pm 0.23$ $2.57\pm 0.01$ $0.54\pm 0.01$ $52.91\pm 0.22$ $2.39\pm 0.01$ $0.52\pm 0.01$ $53.09\pm 0.14$ $1.77\pm 0.02$ $0.53\pm 0.01$ $53.17\pm 0.19$ $3.44\pm 0.02$ $0.52\pm 0.01$
@@ V-A2 Model Setting
In this work, the lag window size $\tau$ is searched over $[7,27]$, the constant $k$ is searched from $[0.02,2]$, and the sparsification threshold $s$ is searched from $[0.25,0.85]$. The number of stock indicators $F$ is set to 4. The number of heads $h$ is searched over $[2,30]$, and the number of layers $L$ is searched from $[2,6]$. In addition, we adopt the PReLU as the activation function, and the AdamW as the optimizer. The learning rate is searched from $10^{-4},2\times 10^{-3}]$ weight decay is searched from $10^{-4},10^{-3}]$, and training epoch is set to 800. In this work, the number of trends $\alpha$ is set to 2, and forecast step $\phi$ is set to 1 following previous works [37, 9, 54, 55], meaning a classification task of predicting next-day stock movements.